in a new system the unit of mass is made 10 times, unit of length is made 1/100 times and unit of time is made 10 times. The magnitude of 1J  in the new system of units is 10^x. find x

Grade:11

3 Answers

Mrinal Kumar
8 years ago
0
Hasan Naqvi
8 years ago
1 J = 1 kg m2/s2 
 
 
1 kg = 0.1 nmu (new mass unit)
1m = 100 nlu (new length unit)
 
thus magnitude of of 1J = 0.1 * 1002= 103
 
x=3
Hasan Naqvi
8 years ago
Sorry didn’t notice the entire question.
 
1 J = 1 kg m2/s2 
 
1 kg = 0.1 nmu (new mass unit)
1m = 100 nlu (new length unit)
1s=  0.1 ntu (new time unit)
thus magnitude of of 1J = 0.1 * 1002/ 0.12= 105
 
x=5